OpenStudy (anonymous):

Solve the equation. Round the result to the nearest hundredth. -15x + 73 = 26

OpenStudy (anonymous):

73-26=15 47=15x x=47/15 x=3.133333333 (round to the nearest hudredth) x=3.13 right?? example.. Rounding Decimals to the Nearest Hundredth Rounding decimals is very similar to rounding other numbers. If the thousandths place of a decimal is four or less, it is dropped and the hundredths place does not change. For example, rounding 0.843 to the nearest hundredth would give 0.84. If the thousandths place is five through nine, the hundredths place is increased by one. The decimal 0.846 rounded to the nearest hundredth is 0.85. correct me if im wrong

OpenStudy (anonymous):

okay thanks thats what I got

OpenStudy (anonymous):

ok :)